3. Renting a BIXI bike in Montreal costs various prices depending on the time frame. Mr. Delaney is interested in renting a bike for the 3-day Osheaga festival in July. The rate is $5 for a 24-hour period and $1.75 for every 30 additional minutes following that.





 a) Write a linear equation to model the cost, C, of the rental per hour, h.





 b) How much will Mr. Delaney be charged if he has the bike for 27 hours?


1
Expert's answer
2021-04-15T07:33:36-0400

a.) Linear equation

"Cost=5+(1.75\times2)h"

"Cost=5+3.5 h"

Where 5 is fixed cost, 3.5 is change in unit cost and h is number of hours variation.

b.) Total charged

Total hours 27.

"Extra hours=27-24=3"

"Cost=5+(3.5 \times3)=15.5"

Cost=$ 15.5